OpenStudy (anonymous):

Find the perimeter of the square. A. 13.0 cm B. 24.0 cm C. 26.0 cm D. 42.25 cm

TheSmartOne (thesmartone):

Where is the square?

TheSmartOne (thesmartone):

@calebwill123

OpenStudy (anonymous):

it has 6.5 all around it

OpenStudy (anonymous):

the square

OpenStudy (anonymous):

do you know it

OpenStudy (anonymous):

well

TheSmartOne (thesmartone):

Perimeter for a square is 4 times the sides

TheSmartOne (thesmartone):

so in your case 4x6.5=26

OpenStudy (anonymous):

ok im just in a rush

TheSmartOne (thesmartone):

So C is your answer.

OpenStudy (anonymous):

so c

TheSmartOne (thesmartone):

Exactly

OpenStudy (anonymous):

yea thx

OpenStudy (anonymous):

Find the perimeter of a rectangle with l = 3.2 m and w = 0.5 m. A. 1.6 m B. 3.2 m C. 3.7 m D. 7.4 m

TheSmartOne (thesmartone):

Perimeter of a rectangle = 2 (L+W)=2(3.2+.5)=2(3.7)=7.4

OpenStudy (anonymous):

so the ANSWER IS srry caps

TheSmartOne (thesmartone):

D.

OpenStudy (anonymous):

ok thx dont go yet

TheSmartOne (thesmartone):

ok

OpenStudy (anonymous):

What is the perimeter of the polygon? A. 56 cm B. 84 cm C. 112 cm D. 168 cm

OpenStudy (anonymous):

24 25 and 7

OpenStudy (anonymous):

r the sides

OpenStudy (anonymous):

do you know it

OpenStudy (anonymous):

@J-bird @laylalove14

OpenStudy (anonymous):

@sumanta

OpenStudy (anonymous):

anyone know it

OpenStudy (anonymous):

@TheSmartOne

TheSmartOne (thesmartone):

add all the sides together to get the perimeter 24+25+7=56

OpenStudy (anonymous):

so 76 thx

OpenStudy (anonymous):

no

OpenStudy (anonymous):

not 76

TheSmartOne (thesmartone):

So 56 is your answer.

TheSmartOne (thesmartone):

A....

OpenStudy (anonymous):

What is the perimeter of the polygon? A. 101.9 cm B. 102.1 cm C. 102.4 cm D. 102.7 cm

OpenStudy (anonymous):

26 15.8 37.6 23.3 r thes sides

OpenStudy (anonymous):

@smartgirl14

TheSmartOne (thesmartone):

add all those numbers and you get the perimeter.

OpenStudy (anonymous):

plz tell me im in a rush

OpenStudy (anonymous):

i got two more so plz then i gotta go

OpenStudy (anonymous):

one more act

OpenStudy (anonymous):

@love_to_love_you

OpenStudy (anonymous):

someone plz tel me

OpenStudy (anonymous):

@J-bird help me

OpenStudy (anonymous):

you know it

OpenStudy (anonymous):

@jose12345 @jakewoln @J-bird

OpenStudy (anonymous):

anyone know it

OpenStudy (anonymous):

anyone know it

OpenStudy (anonymous):

helloooo

OpenStudy (anonymous):

@TheSmartOne @thunder44554431

TheSmartOne (thesmartone):

Just add all those numbers. Don't you have a calculator...

OpenStudy (anonymous):

no i dont

TheSmartOne (thesmartone):

102.7 is your answer

OpenStudy (anonymous):

i do but dont have it with me lost it

OpenStudy (anonymous):

thx dude big help

OpenStudy (anonymous):

one more

OpenStudy (anonymous):

Which equation could be used to find the perimeter of a rectangle that has a width of 18 ft and a length of 24 ft? A. P = 18 · 24 B. P = 2 · 18 + 24 C. P = 2 · 18 + 2 · 24 D. P = 16 + 24

OpenStudy (anonymous):

Which equation could be used to find the perimeter of a rectangle that has a width of 18 ft and a length of 24 ft? A. P = 18 · 24 B. P = 2 · 18 + 24 C. P = 2 · 18 + 2 · 24 D. P = 16 + 24

OpenStudy (anonymous):

Which equation could be used to find the perimeter of a rectangle that has a width of 18 ft and a length of 24 ft? A. P = 18 · 24 B. P = 2 · 18 + 24 C. P = 2 · 18 + 2 · 24 D. P = 16 + 24

OpenStudy (anonymous):

omg

OpenStudy (anonymous):

Which equation could be used to find the perimeter of a rectangle that has a width of 18 ft and a length of 24 ft? A. P = 18 · 24 B. P = 2 · 18 + 24 C. P = 2 · 18 + 2 · 24 D. P = 16 + 24

TheSmartOne (thesmartone):

Which equation could be used to find the perimeter of a rectangle that has a width of 18 ft and a length of 24 ft? C. P = 2 · 18 + 2 · 24

OpenStudy (anonymous):

Which equation could be used to find the perimeter of a rectangle that has a width of 18 ft and a length of 24 ft? A. P = 18 · 24 B. P = 2 · 18 + 24 C. P = 2 · 18 + 2 · 24 D. P = 16 + 24

TheSmartOne (thesmartone):

Why are you posting it like 5 times?

TheSmartOne (thesmartone):

C is the correct answer.

OpenStudy (anonymous):

bec the ?????????????????????

OpenStudy (anonymous):

Which equation could be used to find the perimeter of a rectangle that has a width of 18 ft and a length of 24 ft? A. P = 18 · 24 B. P = 2 · 18 + 24 C. P = 2 · 18 + 2 · 24 D. P = 16 + 24

OpenStudy (anonymous):

see it

OpenStudy (anonymous):

do you know it

TheSmartOne (thesmartone):

2(l+w)= 2l+2w

OpenStudy (anonymous):

c

TheSmartOne (thesmartone):

THats what is said...

OpenStudy (anonymous):

ok thx big help

OpenStudy (anonymous):

i made a 100 thx dude thx thx thxxxxxxxxxxxx

TheSmartOne (thesmartone):

i*

TheSmartOne (thesmartone):

no problem..

OpenStudy (anonymous):

ok bye fan me

TheSmartOne (thesmartone):

Why should I fan you?

TheSmartOne (thesmartone):

@calebwill123

TheSmartOne (thesmartone):

I am the one who helped you.. so you should be fanning me..